# Diverging network architecture of the C. elegans connectome and signaling network

### v1

@@ Results
We found statistical evidence supporting the hypothesis that the signaling network also displays the defining feature of rich-club behavior: neurons with a degree of $k$ or higher highly connected to other neurons of degree $k$ or higher (Fig. 5 a). The signaling network exhibited a “dual rich club” with two clear tiers: the first tier contained 16 neurons of degree 25 or higher, whereas the second and more exclusive tier contained five neurons of degree 40 or higher (Fig. 5 a,b).
@@
Surprisingly, the sets of neurons that comprised the rich clubs in the signaling network were nearly disjoint when compared to the rich club in the anatomical network (Fig. 5 b,c). Neurons that exhibited the highest degree in the anatomical network tended to have lower degrees in the signaling network (Pearson correlation $r=0.22$) (Fig. 5 b red shading) while neurons with the greatest degree in the signaling network were less well-connected in the anatomical network (Fig.  5 b blue and green shading). The only exceptions were the two neurons present in both rich-clubs: command interneurons AVEL and AVER. That this bilateral pair of neurons is highly connected in both networks is consistent with its prominent role in the locomotory circuit driving backwards locomotion [59, 61, 60, 63].
@@
We next asked: how are the signaling network’s rich-club neurons distributed across its communities? Do they span multiple communities, linking them to one another, or do they predominantly congregate in one community? Here we consider both tiers of the rich club together ($k>25$). Community VI was highly enriched for hub neurons, containing 10 out of the 16 hub neurons including AVEL and AVER. Rich club neurons also make up a large fraction of this community, approximately 21%. Given that rich club neurons are, by definition, interconnected, it is then consistent that community VI is also the most highly self-interconnected community (Supplementary Fig. 6).
@@
Of the remaining rich club neurons, two were in pharyngeal community I (M3L, M3R), and several were in communities II (OLLR, OLQDR), III (RMDDR) and V (RMDDL). Surprisingly, members of the rich club were fairly equally spread among sensory, inter-, and motor neurons (Supplementary Fig. 9 b). This is in contrast to the anatomical network where the rich club is primarily composed of interneurons (Supplementary Fig. 9 c). These observations suggest that some neurons historically classified as sensory or motor based on their anatomical wiring may also serve a functional role similar to what has conventionally been ascribed to interneurons.
@@
Rich club analysis relies on measures of in- and out-degree. A potential confounding factor is that signal propagation degree is correlated with the number of empirical observations made of each neuron in our experimental dataset (pearsons correlation $r=0.71$, Supplementary Fig.  8 e). We were reassured, however, that rich club members had large degrees well in excess of what we could explain from the correlation (Supplementary Fig. 5 e,f), and not all highly observed neurons are members. Therefore we conclude that membership in the rich club is not a trivial result of being measured often.
@@ Discussion
As we might expect, there are some similarities between the networks of wiring and signaling. Both are complex networks with hallmark small-world organization, i.e., their wirings are not regular nor random, but poised between these two extremes [26]. The correspondence between the community partitions of the two networks, while small, was still larger than chance, as we expect. Strikingly, both networks had a clear community devoted to the pharynx, consistent with its neurobiological role, as it is thought to be a mostly separate unit of neurons in the C. elegans nervous system. This finding suggests that in real-world brains, it is perhaps only the most extremely pronounced network features, such as the delineation of the nearly isolated pharyngeal sub-network, that are preserved across both signaling and anatomical networks. Even then, we find that in the signaling network the pharynx is less isolated than one would expect from the anatomical network. This could be due, in part, to extrasynaptic signaling from pharyngeal neurons such as M3L which have prominent extrasynaptic connections to other parts of the brain [14].
@@
Other than small-worldness, and the community of pharyngeal neurons, nearly all other aspects of the network were found to be different, including the mapping, delineation, and hierarchy of network modules and the composition of each network’s respective rich clubs.
By inspecting the structure of the signaling network, we generate several hypotheses about possible new roles for different classes of neurons.
@@
For example, our analysis finds that the pharynx plays a broadcasting role as the core in “core-periphery” interactions between communities, which suggests that it may send out signals related to pharyngeal pumping and food to the rest of the network. This could help explain how other portions of the brain get access to food intake signals to regulate behaviors such as dwelling [72].
@@
One surprising finding that needs further follow-up is that amphid sensory neurons were found to have more incoming than outgoing edges in the signaling network, suggesting that they may play a potential role as integrators, even though they are primarily known for their role in receiving sensory inputs. Another surprising finding is the diversity of neuron types that are found in the signaling rich club, which span sensory, motor, and interneurons. Together these findings hint at the possibility that the signaling network in C. elegans is much more functionally intertwined than previously thought– neurons are listening to more diverse sets of neural inputs and may take on more diverse roles than previously appreciated.
All these differences between the two networks lead us to believe that the signaling network may facilitate types of neural computation and communication that are not obvious when only evaluating the anatomical network and therefore while there is a correspondence between structure and function it is not a direct mapping.
These differences could be due to many factors. It is not always clear which synaptic outputs are functionally relevant and therefore a neuron having many synaptic outputs does not trivially translate to having many functional outputs [14, 8, 10]. Different assumptions about how signals dissipate as they fan out through the network could also play a role. For example, if a neuron has many anatomical outgoing connections, one could either assume that it has a strong functional impact because it talks to many post-synaptic patterns, or that it has a weak functional impact because its synaptic output is diluted across many paths. In humans, network models that best predict the propagation of focal electrical stimulation take this dilution into account [22] and further work is needed to investigate this in other brains. Additionally, the presence of extrasynaptic signaling invisible to the anatomical network may also explain some of the differences seen in the signaling network [14].
We are reassured to find evidence that the signaling communities are still meaningful, which suggests that the differences may also have meaning. For example, we found that the signaling communities did have alignment to neuron cell roles and class confirming their neurobiological relevance. Most of the communities were enriched for only one of the three neuron classes: sensory, inter-, and motor neurons, as we expected. Additionally, certain communities were enriched for specific neuron functions, as we also expected.
@@
A challenge of any network analysis is that it is subject to the limitations of the experimental measurements used to describe the underlying network. For the signaling network, we are fortunate to be intimately familiar with those limitations. For example, in the calcium imaging experiments used to measure the signaling network, some neurons are observed or stimulated more often than others, and this is quantified [14]. Uneven sampling introduces a bias into which neurons have edges between them. The signal propagation atlas attempts to account for and correct for this bias by only declaring neurons to be functionally connected after employing a statistical framework that explicitly considers both the magnitude of a neurons response as well as the number of observations [14]. Nonetheless, a correlation still persists between the number of times a neuron was observed or stimulated and its in or out degree, respectively, in the signaling network (Supplementary Fig. 8 a, b, e). Here we present evidence to show that this lingering correlation does not solely explain the richness of the members of the anatomical rich club, for example, (Supplementary Fig. 8 e, f) nor can it trivially explain the integrating or broadcasting nature (defined based on the ratio of out to in degree) of the amphid sensory or pharyngeal neurons, respectively (Supplemental Fig. 8 a,b,c,d). In addition, the stochastic block model is implemented in a degree-corrected matter to partition both networks into communities. Future experimental work is needed to more evenly measure signaling across the network to avoid this confound from the start.
@@
Inhibition is likely systematically undercounted in the underlying measurements of the signaling network due to technical limitations of measuring calcium activity. Because of this, were inhibition to be organized in a systematic way in the network, we would likely not capture that in our analysis. Finally, both of the networks considered here are binarized approximations We chose a binarized approach because it avoids additional assumptions that would be needed to extract meaningful and comparable weights for both the anatomical and signaling networks, however, future work could explore whether a weighted network approach yields more similarities.
@@
Network neuroscience has made vast strides in identifying topological properties of neural networks with the goal of understanding the brain. It is often implicitly assumed that a brain’s network properties such as its partition into modules, its rich clubs, and its small-worldness must be properties of the underlying brain itself, and therefore one would not expect them to depend on the precise lens of network description– whether it is anatomical or signaling. Here we directly compare properties of these two descriptions and find that while some properties are conserved in both the anatomical and signaling networks of the nematode C. elegans, many other properties are not. This difference will be important for interpreting network architecture results of newly acquired connectomes [73, 74] and of forthcoming connectomes [75]. We believe efforts to better understand how these different lenses into the brain relate provide an exciting new frontier into network science [76].
@@ Methods
Unlike some common methods for community detection designed to detect “assortative” community structure–e.g. modularity maximization [78] or infomap [79]– the stochastic blockmodel can detect generic types of community structure. These include core-periphery and disassortative motifs. To determine if the detected communities in the signaling network exhibited these types of non-assortative interactions (as well as the more common assortative type), we followed a recently-proposed method [67]. Specifically, we considered community pairs and their within- and between-community densities of connections. Specifically, for communities $r$ and $s$, we consider the within-community densities, $\omega_{rr}$ and $\omega_{ss}$, as well as the between-community density, $\omega_{rs}$. Given these values, we unambiguously classified the interaction between $r$ and $s$ as assortative if $\omega_{rr}>\omega_{rs}$ and $\omega_{ss}>\omega_{rs}$. Conversely, the interaction was disassortative if $\omega_{rs}>\omega_{rr}$ and $\omega_{rs}>\omega_{ss}$. The core-periphery motif corresponded to the case where $\omega_{rr}>\omega_{rs}>\omega_{ss}$. In this case, communities $r$ and $s$ were labeled as the core and the periphery, respectively (note that $\omega_{ss}>\omega_{rs}>\omega_{rr}$ would also be considered a core-periphery motif, but with the roles reversed).
@@
To determine whether the signaling network contains a rich club we calculate the rich club coefficient $\Phi(k)$ across for every degree $k$. The rich club coefficient measures the density of connections among nodes with degree greater than $k$. We then compare $\Phi(k)$ for our network to the average $\Phi(k)_{rand}$ calculated across 1000 random networks. While $\Phi(k)$ was determined to be statistically significantly larger than $\langle\Phi(k)_{rand}\rangle$ from $k=3$ onwards we concentrated on the two local maxima in $\Phi(k)_{norm}$ at $k=25$ and $k=40$.
@@
Extended Data Figure 2: Variability in community assignment by neuron. Pie charts indicate how consistently neurons were assigned to the same community in the structural network. Organized by community. Blue corresponds to the fraction assigned to the consensus community assignment (largest fraction), orange the next most common community assignment then green etc.
@@
[…]

### v2

@@ Results
We found statistical evidence supporting the hypothesis that the signaling network also displays the defining feature of rich-club behavior: neurons with a degree of $k$ or higher, highly connected to other neurons of degree $k$ or higher (Fig. 5 a). The signaling network exhibited a “dual rich club” with two clear tiers: the first tier contained 16 neurons of degree 25 or higher, whereas the second and more exclusive tier contained five neurons of degree 40 or higher (Fig. 5 a,b).
@@
There were two neurons present in both rich clubs: command interneurons AVEL and AVER. This bilateral pair of neurons is very highly connected in both networks, which is consistent with its prominent role in the locomotory circuit driving backwards locomotion [69, 71, 70, 73]. This again illustrates that the most extreme features of the network are preserved across anatomy and signaling. Otherwise, the sets of neurons that comprised the rich clubs in the signaling network surprisingly were nearly disjoint when compared to the rich club in the anatomical network (Fig. 5 b,c). Neurons that exhibited the highest degree in the anatomical network tended to have lower degrees in the signaling network (Pearson correlation $r=0.22$) (Fig. 5 b red shading) while neurons with the greatest degree in the signaling network were less well-connected in the anatomical network (Fig.  5 b blue and green shading). We find that this result persists in both the conservative control and the enforced density controls (Extended Data Fig. 11 d, 12a, & 13 a).
@@
We next asked: how are the signaling network’s rich-club neurons distributed across its communities? Do they span multiple communities, linking them to one another, or do they predominantly congregate in one community? Here we consider both tiers of the rich club together ($k>25$). Community VI was highly enriched for hub neurons, containing 10 out of the 16 hub neurons, including AVEL and AVER. Rich club neurons also make up a large fraction of this community, approximately 21%. Given that rich club neurons are, by definition, interconnected, it is then consistent that community VI is also the most highly self-interconnected community (Extended Data Fig. 7).
@@
Of the remaining rich club neurons, two were in pharyngeal community I (M3L, M3R), and several were in communities II (OLLR, OLQDR), III (RMDDR) and V (RMDDL). Surprisingly, members of the rich club were fairly equally spread among sensory, inter-, and motor neurons (Extended Data Fig. 10 b). This is in contrast to the anatomical network where the rich club is primarily composed of interneurons (Extended Data Fig. 10 c). These observations suggest that some neurons historically classified as sensory or motor based on their anatomical wiring may also serve a functional role similar to what has conventionally been ascribed to interneurons.
@@
Rich club analysis relies on measures of in- and out-degree. A potential confounding factor is that signal propagation degree is correlated with the number of empirical observations made of each neuron in our experimental dataset (pearsons correlation $r=0.71$, Extended Data Fig.  9 e). We were reassured, however, that rich club members had large degrees well in excess of what we could explain from the correlation (Extended Data Fig. 5 e,f), and not all highly observed neurons are members. Therefore we conclude that membership in the rich club is not a trivial result of being measured often.
@@ Discussion
One’s perspective likely influences the degree to which one would expect network features of the anatomical network to be similar to signaling networks. Since they both reflect the same underlying nervous system, it is tempting to assume that key features will be preserved. And indeed, the most obvious features are: the pharynx forms its own community in both anatomy and signaling. In addition, the richest members of the rich club in anatomy are also the richest members of the rich club in signaling.
However, as we consider the dynamical systems nature of the nervous system, it becomes clear that there are fundamental differences between a network’s description of anatomy and a network description of signaling. For example, anatomy describes direct wired interactions, while signaling captures how neural activity travels multi-hop through the network, including feedback, and recurrence. Even superficially basic assumptions about how to interpret nervous system wiring in a network context are unclear. For example, if a neuron has many anatomical outgoing connections, does this suggest a strong signaling impact because it talks to many post-synaptic patterns, or a weak signaling impact because its synaptic output is diluted across many paths? Network models that consider dilution perform well in humans [18] and are one possible avenue for future exploration. Given the challenges, it may not be surprising that network descriptions based on specific assumptions of the underlying dynamics do not necessarily generalize to all brains or even to brains at all [8].
Similarly, as one considers additional biophysical details, it becomes clear that wiring is not the only factor that influences dynamics and that contributions from neuromodulation, extrasynaptic signaling, and gene expression all have important roles to play and are poorly captured by the network description of anatomy alone.
Together, our findings suggest that in real-world brains, it is perhaps only the most pronounced network features, such as the delineation of the nearly isolated pharyngeal sub-network, that we should expect to be preserved across both signaling and anatomical networks.
In both signaling and anatomy, we find that bilaterally symmetric neurons are much more likely to be found in the same communities than across different communities, suggesting a strong role of bilateral symmetry in the brain. How do we interpret the observation that this bilateral symmetry is extremely pronounced in the anatomical wiring (38 times more likely to be in the same community) but less so in signaling (twice as likely)? Here again, fundamental differences in what the networks describe could play a role. Community membership in the anatomical network is influenced entirely by monosynaptic wiring patterns. Therefore, laterality in the wiring patterns one hop away has no bearing on a neuron’s community membership in the anatomical network. By contrast, the signaling network reflects poly-synaptic or multi-hop signaling. Therefore, we would expect laterality in wiring more than one hop away to influence a neuron’s signaling community membership. In this way, we expect the signaling network to be generally more susceptible than the anatomical network to laterality in wiring that occurs anywhere in the network. In addition to this increased sensitivity to laterality in wiring, we also expect the signaling network to reflect laterality in gene expression and other non-wired aspects of the nervous system. Therefore, it is perhaps not surprising that we find that the crisp bilaterality of the connectome is most strongly reflected in the anatomical network and only more weakly reflected in the signaling network.
@@
Beyond the striking examples discussed above, nearly all other aspects of the network were found to be different, including the mapping, delineation, and hierarchy of network modules and the composition of each network’s respective rich clubs.
@@
Despite the challenges of relating the two networks, we note that the network architecture of the signaling network on its own provides intriguing hypotheses about the possible role of neurons and subnetworks in the brain. For example, our analysis finds that the pharynx plays a broadcasting role as the core in “core-periphery” interactions between communities, which suggests that it may send out signals related to pharyngeal pumping and food to the rest of the network. This could help explain how other portions of the brain get access to food intake signals to regulate behaviors such as dwelling [84]. We find that in the signaling network, the pharynx is less isolated than one would expect from the anatomical network, likely due to polysynaptic signaling and also possibly through extrasynaptic signaling of key pharyngeal neurons, such as M3L [27].
@@
One surprising finding is that amphid sensory neurons were found to have more incoming than outgoing edges in the signaling network, suggesting that they may play a potential role as integrators, even though they are primarily known for their role in receiving sensory inputs. Another surprising finding is the diversity of neuron types that are found in the signaling rich club, which span sensory, motor, and interneurons. Together, these findings hint at the possibility that the signaling network in C. elegans is much more functionally intertwined than previously thought– neurons are listening to more diverse sets of neural inputs and may take on more diverse roles than previously appreciated.
@@
A challenge of any network analysis, is that it is only as good as the experimental measurements that describe the underlying network. In general, a caveat of measuring signal-evoked responses is that we are limited by the signal-to-noise ratio of our calcium indicator, and we miss weak, noisy functional connections or those of which we have few measurements. For example, in the calcium imaging experiments used to measure the signaling network, some neurons are observed or stimulated more often than others, and this is quantified [27]. Uneven sampling introduces biases, which neurons have edges between them. The signal propagation atlas attempts to account for and correct for this bias by only declaring neurons to be functionally connected after employing a statistical framework that explicitly considers both the magnitude of a neuron’s response as well as the number of observations [27]. Nonetheless, a correlation still persists between the number of times a neuron was observed or stimulated and its in or out degree, respectively, in the signaling network (Extended Data Fig. 9 a, b, e). Here we present evidence to show that this lingering correlation does not solely explain the richness of the members of the anatomical rich club, for example, (Extended Data Fig. 9 e, f) nor can it trivially explain the integrating or broadcasting nature (defined based on the ratio of out to in degree) of the amphid sensory or pharyngeal neurons, respectively (Extended Data Fig. 9 a,b,c,d). In addition, the stochastic block model is implemented in a degree-corrected manner to partition both networks into communities. We also explore several controls in which we alter the thresholds for edge inclusion. Future experimental work is needed to more evenly measure signaling across the network to avoid this confound from the start.
@@
Inhibition is likely systematically undercounted in the underlying measurements of the signaling network due to technical limitations of measuring calcium activity. Because of this, were inhibition to be organized in a systematic way in the network, we would likely not capture that in our analysis. Finally, both of the networks considered here are binarized approximations. We chose a binarized approach for simplicity, and showed that key results are not sensitive to modest changes in binarization thresholds. However, future work could explore whether a weighted network approach yields more similarities.
A future direction is to try to understand how features of the signaling network may arise from the spatial organization of neurons and synaptic contacts. For example, we expect that extrasynaptic neuropeptide signaling can only occur between pairs of neurons that are sufficiently close so that neuropeptides released from one neuron are sufficiently likely to reach neuropeptide receptors on the other. One challenge is that the locations of peptides and receptors on neurons are not known. More generally, it has been challenging to identify relevant spatial metrics (e.g., distances) between neurons because neurons are extended bodies that often are arranged into dense bundles, and have long processes with extended contact areas across many neighbors [11, 36]. New progress may make this possible in the near future [85].
@@
Here we have directly compared properties of signaling and anatomical descriptions of a brain and find that while some properties are conserved in both the anatomical and signaling networks of the nematode C. elegans, many other properties are not. This difference will be important for interpreting network architecture results of newly acquired connectomes [86, 17] and of forthcoming connectomes [87]. We believe efforts to better understand how these different lenses into the brain relate provide an exciting new frontier into network science [88].
@@ Methods
To determine if the detected communities in the signaling network exhibited these types of non-assortative interactions (as well as the more common assortative type), we followed a recently-proposed method [77]. Specifically, for communities $r$ and $s$, we consider the within-community densities, $\omega_{rr}$ and $\omega_{ss}$, as well as the between-community density, $\omega_{rs}$. Given these values, we unambiguously classified the interaction between $r$ and $s$ as assortative if $\omega_{rr}>\omega_{rs}$ and $\omega_{ss}>\omega_{rs}$. Conversely, the interaction was disassortative if $\omega_{rs}>\omega_{rr}$ and $\omega_{rs}>\omega_{ss}$. The core-periphery motif corresponded to the case where $\omega_{rr}>\omega_{rs}>\omega_{ss}$. In this case, communities $r$ and $s$ were labeled as the core and the periphery, respectively (note that $\omega_{ss}>\omega_{rs}>\omega_{rr}$ would also be considered a core-periphery motif, but with the roles reversed).
@@
To determine whether the signaling network contains a rich club we calculate the rich club coefficient $\Phi(k)$ across for every degree $k$. The rich club coefficient measures the density of connections among nodes with degree greater than $k$. We then compare $\Phi(k)$ for our network to the average $\Phi(k)_{rand}$ calculated across 1000 random networks in which the edges were randomly rewired while preserving the original degree of each node (using graph_tool.generation.random_rewire with the model ‘configuration’). While $\Phi(k)$ was determined to be statistically significantly larger than $\langle\Phi(k)_{rand}\rangle$ from $k=3$ onwards we concentrated on the two local maxima in $\Phi(k)_{norm}$ at $k=25$ and $k=40$.
Network Thresholding Controls
We changed the thresholds of the networks to perform two different controls.
We performed a comparison of the two networks, where both networks were thresholded to be more conservative in terms of their inclusion criteria for connections. For the signaling network, we used a threshold of $q=0.01$ instead of $q=0.05$, leading to a density of $d_{s}=0.021$ (instead of $0.033$). To include a pair of neurons in the anatomical network, we required that they must have at least three synaptic contacts between them in at least one of the connectomes. This led to a density of $d_{a}=0.028$ (instead of $0.092$).
Additionally, in Extended Data Figures 12 and 13, we compared networks while enforcing their densities to be the same. We did so by thresholding the networks in either of two ways. For the anatomical network in Extended Data Figure 12, we required that for a pair of neurons to be anatomically connected the average synapses across the four connectome datasets be at least 1.25 synapses (instead of greater than zero), since this corresponds to a network connection density of $d_{a}=0.033$ (comparable to the connection density of the signaling network $d_{a}=0.033$). We then compared this thresholded anatomical network to the original signaling network. We separately performed the converse adjustments in Extended Data Figure 13. There, we required that the signaling network have a q-value threshold of $q<=0.18$ (instead of $q<=0.05$) since this corresponds to a network connection density of $d_{s}=0.091$ (comparable to the connection density of the anatomical network $d_{s}=0.092$). We then compared this thresholded signaling network to the original anatomical network.
@@
Network
Original
Signaling
Original
Anatomy
Conservative
Signaling
Conservative
Anatomy
Enforced density
Signaling
Enforced density
Anatomy
L/R neurons
same-community
odds
2 38 1.3 8 2.43 78
Extended Data Table 1: How much more likely L/R pairs of neurons are to be in the same community than not depending on network thresholds.
Extended Data Figure 2: Variability in community assignment by neuron. Pie charts indicate how consistently neurons were assigned to the same community in the signaling network, organized by community. Blue corresponds to the fraction assigned to the consensus community assignment (largest fraction), orange the next most common community assignment then green etc.
@@
Extended Data Figure 4: Overlap between anatomical and signaling networks was not improved by forcing the number of communities to be the same. a) Overlap between the anatomical and signaling communities. When partitioning the networks via stochastic block modeling, the anatomical network was forced to be partitioned into the same number of communities as the optimal partition for the signaling network (6 communities). b) Relative enrichment of overlap between anatomy and signaling reported as a z-score for each anatomy-signaling community pair (* $p<0.05$, ** $p<0.01$, p-values from the same distribution as z-score, FDR-adjusted). c,d) Same as a and b however, with the partitioning for the signaling network forced to be the same number of communities as the optimal partition for the anatomical network (20 communities). Anatomical communities in all sub-panels are organized by best-matching via a similarity matrix and the Hungarian method.
@@
Extended Data Figure 6: Anatomy is net feed-forward along the sensory to inter- to motor neuron pathway while signaling is ambiguous. a) Number of edges between sensory, inter-, and motor neurons for anatomy (top) and signaling (top). Arrow thicknessis proportional to the number of edges along the arrow. b) log of the ratio of the number of edges forwards to backwards along the sensory $\rightarrow$ inter $\rightarrow$ motor pathway for each possible leg of the pathway: sensory $\rightarrow$ interneurons, inter- $\rightarrow$ motor neurons and sensory $\rightarrow$ motor neurons. Anything above 0 indicates more forwards than backward edges implying feed-forwardness, anything below indicates more backward than forward edges implying feed-backwardness. Red indicates the true log of the ratio of the number of edges. Grey violin plots are derived from a control in which assignments of the labels "sensory", "inter" and "motor" are shuffled. c) Same as b) except the grey violin plots are derived instead from a control where the network is shuffled in a way that preserves both the total in and out-degree of each neuron. (* p<0.05, ** p<0.01, p-values calculated via a two-sided t-test)
@@
Extended Data Figure 11: A more conservative thresholding of both networks yields similar results. a) Enrichment of thresholded anatomical and signaling communities for neuron role (* $p<0.05$, ** $p<0.01$, p-values FDR-adjusted). b) The anatomical rich club is still non-overlapping with the signaling rich club, other than neurons AVEL/R. c) Overlap (counts) between membership of thresholded anatomical and signal propagation communities. d) Relative enrichment of overlap is reported as a z-score for each thresholded anatomy-signaling community pair (* $p<0.05$, ** $p<0.01$, p-values from same distribution as z-score, FDR-adjusted).
Extended Data Figure 12: Adjusting the threshold of the anatomical connectome to match the density of the signaling network yields similar results. The anatomical network was thresholded based on the number of synapses needed for there to be an edge between two neurons to closely match the density of the signaling network $d_{a}=0.033$. The thresholded anatomical network was optimally partitioned into 16 communities. a) The anatomical rich club is still non-overlapping with the signaling rich club, other than neurons AVEL/R. The anatomical rich club consists mostly of the same neurons as originally (9 out of the 15 original neurons are the same, with 4 new neurons). b) Enrichment of anatomical communities for neuron role (* $p<0.05$, ** $p<0.01$, p-values FDR-adjusted). c) Overlap (counts) between membership of anatomical and signal propagation communities. d) Relative enrichment of overlap is reported as a z-score for each anatomy-signaling community pair (* $p<0.05$, ** $p<0.01$, p-values from same distribution as z-score, FDR-adjusted).
Extended Data Figure 13: Adjusting the threshold of the signaling network to match the density of the anatomical network yields similar results. The signaling network was thresholded based on the maximum q-value of a connection between two neurons needed for there to be an edge between two neurons to closely match the density of the anatomical network $d_{s}=0.091$. The thresholded signaling network was also optimally partitioned into 6 communities. a) The signaling rich club is still non-overlapping with the anatomical rich club, other than neurons AVEL/R. All neurons in the signaling rich club were also a part of the original rich club. b) Enrichment of signaling communities for neuron role (* $p<0.05$, ** $p<0.01$, p-values FDR-adjusted). c) Overlap (counts) between membership of anatomical and signal propagation communities. d) Relative enrichment of overlap is reported as a z-score for each anatomy-signaling community pair (* $p<0.05$, ** $p<0.01$, p-values from same distribution as z-score, FDR-adjusted).
Extended Data Figure 14: Modular structure at different thresholds. Circular dendograms showing community assignments at different hierarchical levels for anatomy (left) and signaling (right) at the different binarization thresholds used in the paper and in the controls.